# On the foundation of quantum decision theory

### v2

@@ 2.5 A survey of quantum probabilities
These results give an extended interpretation of the operator $A=A^{\theta}$: There is a simple formula for all expectations in terms of the operator. On the other hand, the set of such expectations determine the state of the system. Also on the other hand: If $A$ is specialized to the operator of an indicator function, we get back Born’s formula, so the consequences are equivalent to this formula.
A consequence of 3) above is that $\theta=\theta^{b}$ does not need to be maximal in order that a Born formula should be valid. The version 2) of Born’s formula is also valid when the basic variable $\theta^{a}$ behind the densitity operator $\rho$ is not maximal, under a certain assumption: Assume that $\theta^{a}=f(\lambda^{a})$, where $\lambda^{a}$ is maximal and the conditional probability distribution of $\lambda^{a}$, given $\theta^{a}$ is uniform.
Measurements of empirical variables are discussed in Helland (2021), but here we will look at the simpler case of a perfect measurement. Assume that we know the state $|\psi\rangle$ of a system, and that we want to measure a new theoretical variable $\theta^{b}$. This can be discussed by means of the projection operators $\Pi_{j}^{b}=|b;j\rangle\langle b;j|$. First observe that by a simple calculation from Born’s formula
@@
It is interesting that Shrapnel et al. (2017) recently simultaneously derived both the Born rule and the well-known collapse rule from a knowledge-based perspective. The collapse rule is further discussed in Helland (2021), but in this article we will just assume this derivation as given. Then, after a perfect measurement $\theta^{b}=u_{j}^{b}$ has been obtained, the state changes to
@@ 2.6 Born’s rule in a decision setting
Go back to the person $C$ who is facing at least two related maximal decisions, with corresponding decision variables $\theta$ and $\eta$. Each of these decisions may be composed of several partial decisions; then the decision variables may be thought of as vectors. However, we will make the simplifying assumption that each partial decision has a finite number of outcomes, say possible actions; then the total decision also has a finite number of outcomes, and without loss of generality, we can order these actions as $\{a_{i};i=1,2,..,r\}$.
Acording to Kahnemam (2011), some of our decisions are slow, and sometimes, in fact in very many cases, we have to make fast decisions, relying on earlier experience, perhaps in a subconscious way, relying on some abstract ideals that we may have in our minds at the moment of decision, or, alternatively, ideals that have determined our earlier decisions. This is consistent with the model assumptions that we have made above in connection to Born’s rule. And the fact that such fast decisions may be modelled by quantum probabilities, is consistent with a lot of empirical findings, see Pothos and Busemeyer (2022).
Focus again on the person $C$ and his fast decisions. Depending upon the situation that he is in, he may just have done several irrelevant decisions, collected in the variable $\theta$, but at the same moment, he is trying to focus on the problem that he really is interested in, and the corresponding (maximal) variable is called $\eta$. There is not much lack of generality in assuming that these two decision processes have the same number of actions $r$; we can only add some abstract irrelevant actions to one of the decision variables.
According to my interpretation of quantum theory and of decision making, each of the two decision variables has an operator connected to it, and the eigenspaces of such an operator correspond to a concrete outcome of the decision process. The probabilities are calculated as in (3). The application of these theoretical considerations is best studied by a simple example.
@@ 3 Quantum decision theory in practice; an example
Consider a situation with one doctor and one patient. The doctor has the choice between two mutually excluding medicines $A$ and $B$.
The doctor starts by asking a number of question, and he obtains answers. Result: A state $|\psi\rangle$ for the patient. This is the starting point of my model of what is going on in the mind of the doctor during the decision process. The state is modelled by some vector in an underlying Hilbert space, connected to the mind of the doctor.
@@
However, from our point of view, the most interesting case is where there is some coupling between the two medicines. The doctor must make fast decisions. In his mind he may have a lot of other decisions that he has made or is going to make. As a model, these may be collected in a decision variable $\theta$. But now he focuses on the actual patient. The decisions that he is going to make on this patient, are modelled, and collected in the decision variable $\eta$. In the doctor’s mind at the moment in question, is also all the information that he posesses about the medicines $A$ and $B$. As a doctor, he has ideals learned through many years of education and experience; these ideals can be modelled in terms of some abstract perfectly rational being $D$.
We will propose a quantum model with Hilbert space $\mathcal{H}$. The projector upon the subspace indicating that $A$ helps, is denoted by $\Pi_{A}$.
@@
To continue the model, consider two orthogonal spaces in $\mathcal{H}$: $V_{B}$: Indicating that $B$ helps. The projector upon this space is $\Pi_{B}$. And $V_{B}^{\mathrm{perp}}$: Indicating that $B$ does not help. The projector upon this space is $\Gamma_{B}=I-\Pi_{B}$.
@@ 3.1 Some consequences
For a tentatively illustration, see Figure 1.
@@
Finally, note that one my have
@@
All these results are well known properties of quantum probabilities; these probabilities simply do not follow Kolmogorov’s laws.They are also well known properties in various empirical examples given in the literature of Quantum Decision Theory. My task now will be to relate the quantum probabilities, as observed in practice, to the assumptions made earlier in this article.
4 Discussion
4.1 A summary of assumptions
It is crucial that this article relies upon my earlier work on the foundation of quantum theory. This work in turn relies on several assumptions, as stated here in connection to the discussion around Theorem 1 in Subsection 2.1:
1) The variables, here decision variables are divided into accessible and inaccessible ones. A decision is said to be accessible to a person $C$ if and only if $C$ is able to carry out the decision.
2) All accessible decision variables can be seen as functions of some large inaccessible variable. In practice this may be taken to mean that all earlier experiences of $C$, among other things related to his earlier decisions, are collected in an hypothetical large variable $\phi$. Note that $C$ has not grown up in isolation. He is the result of interaction with other people, in particular the result of a specific human culture. Depending upon circumstances, he may have made specific conscious decisions to free himself from this culture, but nevertheless, the background is there.
3) There exist maximal accessible decision variables. As stated earlier, these variables are connected to really difficult decisions. Such decisions may be conscious or unconscious. I will see it as fruitful to distinguish between what Kahneman (2011) calls fast and slow decisions. Many of our decisions have to be fast, and I see the theory of this article as specially connected to these fast decisions.
4) To lead to a specific quantum theory, we seem to have to look at situations where there are two different, mutually exclusive decision processes in the mind of $C$ such that both can be seen as maximal. He or she then has to make a choice: Focus on the decision process connected to the decision variable $\theta$ or focus on the decision process connected to the decision variable $\eta$. Again, this choice may be conscious or unconscious. I will claim that at least in some cases, it is a fast decision, coupled solely to the unconsciousness of $C$.
In addition to these assumptions, this article also relies on specific assumptions behind the Born rule, the means for calculating quantum probabilities -see Subsections 2.4-2.6:
5) The person with whom the probabilities are associated, believes in the statistical likelihood principle. Most statisticians believe in this principle, it is connected to the socalled Birnbaum’s theorem: It can be deduced from simpler, more inuitive principles, see Helland (2021).
6) Both $C$ and the person with whom the probabilities are associated have ideals, ideals that can be modelled by some abstract being $D$, which is seen by the actors to be perfectly rational, as made explicit by the Dutch Book Principle. For different persons, these ideals may be very different. They may be connected to other, concrete persons, or they may just be abstract. This point is very crucial for me. We all go through life having certain ideals, and in most cases, this has a very positive effect on our lives.
But the ideals may also be connected to persons that I myself have little sympathy for. Extreme cases exist, like Adolf Hitler in Germany before and under the second world war, or Mafia bosses behind criminal decisions made by gang members. Extending this to my own opinions, I am also tempted to point at Vlademir Putin behind certain decisions made by 70 percent of the Russian population, and Donald Trump behind many of the decisions made by rightwing Americans.
What are then my own ideals? During my work with the quantum foundation, I have had a strong religious faith, as discussed explicitly in the two articles Helland (2022d) and Helland (2023d). I do believe in the existence of a God, and I have sympathies with the best sides of all religious beliefs. Personally, I have chosen to be a Christain, but I do well understand people that have made other choices here. Religion is definitely connected to culture.
In making concrete decisions, I also rely on various ideals related to specific persons that I look up to.
4.2 Decisions, learning, and cultures
As stated in the Introduction, there is a close connection between decisions and learning. The learning processes start when we are infants, and continues throughout our lifes. The decisions that we make in our daily life can be seen as a product of this.
That implies that the decisions that we make today, partly may be a product of the interactions that we thoughout life have made with other people. In a somewhat larger perspective, the sum of all this may be called culture. The author Ralph B Stacey has written in that connection: ‘Culture is a set of attitudes, opinions and convictions that a group of people share, on how one shall behave, how things shall be evaluated and be done, what questions are important and what answers that can be accepted. The most important elements in a culture are unconscious, and cannot be changed from the outside.’
It should be emphasized that belonging to a culture has many positive sides, but a culture may become highly negative if it leads to hate and mistrust of people from other cultures.
It should be mentioned in this connection that Helland (2034b) is an attempt to create common results based on three different scientific cultures. I see science as a very important background for several of our decisions, but I see it as negative that, partly, science may be divided into disjoint cultures, with too few attempts of explicit communication between the different cultures.
4.3 Everyday decisions
Most everyday decisions are not related to a maximal decision variable. But sometimes a decision requires the most of our resources, say in certain situations during driving a car. Note that my discussion of Quantum Decision Theory, in particular the discussion of Born’s rule, is not only valid when the variables are maximal. This is an important point, and it is perhaps not properly emphasized in the discussion above. Again I refer to Mogiliansky et al. (2009) for theoretical and practical implications.
This should imply that Quantum Decision Theory applies to many everyday decisions, and this is confirmed empirically by the QDT literature.
What Born’s rule requires, is some observer to whom the probabilities are connected, a person that that has a mental model that determines these probabilities. In the QDT literature, this is represented by the researcher that carries out the relevant investigation.
The exception to this use of Born’s rule in a decision context, may be what Kahneman (2011) calls decisions based upon slow thinking, decision where the actor $C$ actively uses his conscious self in the process of making the decisions. It is very unlikely that such decisions can be modeled by QDT.
4.4 Decisions made by political leaders
Political leaders have to make decisions that may have consequences for many people. This gives them a great responsibility. Most democratic leaders are able to live up to this responsibility, but there are exceptions.
These responsibilities are spread across a large number of concrete issues. One important issue is the climate question. It is a hope that political leaders in all countries now may be able to make the right decisions here.
Many of the issues that political leaders have to make decision on, are really difficult, often too difficult to be left to a single person. In democratic countries, joint decisions are often made by a group of communication persons. Such joint discussions are not discussed in the present article, but they could have been, in the light of a version of Theorem 1 which is valid for a group of communicating actors, see Helland (2024a).
It is a hope that the really important decisions made by political leaders are based on Kahneman’s slow thinking, conscious decisions that are outside the scope of this article.
Certain political leaders may make decisions that really have cruel consequences. Some examples are 1) the cruel attack on Ukraine initiated by Vlademir Putin in February 2022; 2) the attack on innocent sivilians ordered by the leaders of Hamas in October 2023; 3) the continuous attack on innocent people in Gaza lead by Israel’s leaders.
One constant threat to humanity are the nuclear weapons, weapons that now are spread across too many countries. Any use of nuclear weapons will lead to enormous sufferings, and may in the worst case be the start of a process that can lead to the extinction of the human race.
One clear wish is that they never should be used again. How can that be avoided? My idea here is that this may be connected to the observation that all people, also political leaders may have some religious feelings, at least some thoughts around their destiny after death. May be the following thought could be spread and made common: The political leader who ever dares to order some atomic bomb to be released, deserves an extremely hard punishment after death. If this could be believed by all our leaders, the atomic arsenal spread around the world will be totally useless, and negotiations around the reduction of this arsenal might start. This is of course an idealistic thought, but on this important issue, one should really try to think in a new way.
@@ 5 Concluding remarks
Quantum decision theory seems to be applicable in many cases where decisions are to be made, primarily in the cases which Kahneman (2011) calls decisions of type 1, decisions made through fast thinking. When a prior is available, an alternative is a Bayesian decision theory. This theory will follow the Kolmogorov rules, and apart from the existence of a prior, the theory does not require any special assumptions.
In this article I have concentrated on decisions made by a single person. But the theory can be extended to joint decisions made by a group of communicating persons. Then these persons must have common ideals, as modelled by a perfectly rational higher being $D$. The ideals must be common in the setting determined by the relevant decision variables.
@@
As a basis for quantum decision theory as discussed here, lies the idea of two complementary decision variables at the same time in the mind of some person or in the joint minds of a group of communicating persons. I repeat that two variables are here called complementary if they are really different (not in one-to-one correspondence) and maximal as accessible variables. Another essential assumption is the existence of ideals that are perceived as perfectly rational, modelled by a hypothetical ‘ideal higher being’.
The quantum probabilities may be though of as relative to this ‘ideal higher being’, hence they must be looked upon as absolute probabilities, not conditioned in any other way than by our past knowledge. This is confirmed by empirical findings in the quantum decision literature.
There are cases where powerful political leaders may be roughly modelled as such ‘ideal higher beings’, relative to the influence they have on other people and their decisions. A fundamental requirement should then be that the leaders obey certain basic humanitarian conditions, including respect for other people’s life. Unfortunately, this goal is not always satisfied in practice.
@@
The assumptions may in certain situations be satisfied by very many different pairs of decision variables, perhaps leading to mental difficulties in the first case and to partly complementary world views in the last two cases. Then in these cases it might imply deep conflicts between those persons or groups if some necessary conditions are there. We can see many examples of this phenomenon in the world today.
It is a hope that this article may lead to an increased focus on making the right decisions, the best for a happy and meaningful private life, and in a global perspective, the best decisions in order to obtain peace on this earth and a viable future for all people.
6 Aknowledgements
@@
Refererences
Aerts, D. (2009). Quantum structure in cognition Journal of Mathematical Psychology 53, 314-348.
Allais, M. (1953). La psychologie de l’homme rationnel devant le risque: Critique des postulats et axiomes de l’ecole Ameéricaine. Econometrica 21, 515-550.
Ashtiani, M. and Azgomi, M.A. (2015). A survey of quantum-like approaches to decision making and cognition. Mathematical Social Sciences 75, 49-80.
Auffeves, A. and Granger, P. (2019). Deriving Born’s rule from an inference to the best explanation. arXiv:1910.13738 [quanr-ph].
Busemeyer, J.R. and Bruza (2012). Quantum Models of Cognition and Decision. Cambridge University Press, Cambridge.
Busch, P. (2003). Quantum states and generalized observables: A simple proof of Gleason’s Theorem. Physical Review Letters, 91(12), 120403.
Busch, P., Lahti, P., Pellonpää, J.-P., & Ylinen, K. (2016). Quantum measurement. Berlin: Springer.
Campanella, M., Jou, D., & Mongiovi, M. S. (2020). Interpretative Aspects of Quantum Mechanics. Matteo Campella’s Mathematical Studies. Cham, Switzerland: Springer.
Caves, C.M., Fuchs, C.A., Manne, K., and Renes, J.M. (2004). Gleason-type derivations of the quantum probability rule for generalized measurements. Foundations of Physics 34, 193-209.
Ellsberg, D. (1961). Risk, ambiguity and the Savage axioms. Quaterly Journal of Economics. 75, 643-679.
Friston, K.J., Daunizeau, J. and Kiebel, S.J. (2009). Reinforcement learning or active inference? PLoS ONE 4 (7) e6421.
Helland, I.S. (2021). Epistemic Processes. A Basis for Statistics and Quantum Theory. Springer, Cham, Switzerland.
Helland, I.S. (2022a). On the diversity and similarity of mathematical models in science. American Review of Mathematics and Statistics 10 (1), 1-10.
Helland, I.S. (2022b). On reconstructing quantum theory from two related maximal conceptual variables. International Journal of Theoretical Physics 61, 69. Correction (2023) 62, 51.
Helland, I.S. (2022c). The Bell experiment and the limitations of actors. Foundations of Physics 52, 55.
Helland; I.S. (2022d). On religious faith, Christianity, and the foundation of quantum mechanics. European Journal of Theology and Philosophy 2 (1), 10-17.
Helland, I.S. (2023a). On the Foundation of Quantum Theory. The Relevant Articles. Eliva Press. Chisinau, Moldova.
Helland, I.S. (2023b). On the Bell experiment and quantum foundation. Journal of Modern and Applied Physics 6 (2), 1-5.
Helland, I.S. (2023c). Possible connections between relativity theory and a version of quantum theory based upon theoretical variables. arXiv: 2305.15435 [physics.hist-ph].
Helland, I.S. (2023d). Quantum mechanics as a theory that is consistent with the existence of God. Dialogo Conferences and Journal 10 (1), 127.134.
Helland, I.S. (2024a). An alternative foundation of quantum mechanics. arXiv: 2305.06727 [quant-ph]. Foundations of Physics 54, 3.
Helland, I.S. (2024b). Towards optimal linear prediction. Submitted.
Kahneman, D. (2011). Thinking, Fast and Slow Penguin Books. London.
Kahneman, D., Sibony, O. and Sunstein, C.R. (2021). Noise. A Flaw in Human Judgement. Brockman, Inc., New York.
Mogiliansky, A.L., Zamir, S. and Zwirn, H. (2009). Type indeterminacy: A model of the KT(Kahneman-Tverski)-man. Journal of Mathematical Psychology 53 (5), 349-361.
Newell, B.R., Lagnado, D.A. and Shanks, D.R. (2015). Straight Choices. The Psychology of Decision Making. 2. edition. Psychology Press, London.
Pothos, E.M. and Busemeyer, J.R. (2022). Quantum cognition. Annual Review of Psychology 73, 749-778.
Shrapnel; S., Costa, F. & Milburn, G. (2017). Updating the Born rule. arXiv: 1702.01845v1 [quant-ph].
von Neumann, J. (1927). Wahrscheinlichkeitstheoretischer Aufbau der Quantenmechanik. Nachrichten von der Gesellschaft der Wissenschaften zu Göttingen, Mathematisch-Physikalische Klasse, 1927, 245–272.
Wright, V.J. and Weigert, S. (2019). A Gleason-type theorem for qubits based on mixtures of projective measurements. journal of Physics A: Mathematical and Theoretical 52 (5), 055301.
Wright, V.J. and Weigert, S. (2021). General probabilistic theories with a Gleason-type theorem. Quantum 5, 588.
Yukalov, V.I. and Sornette, D. (2014). How brains make decisions. Springer Proceedings in Physics 150, 37-53.
Zwirn, H. (2016). The measurement problem: Decoherence and convivial solipsism. Foundations of Physics 46, 635-667.
Zwirn, H. (2020). Nonlocality versus modified realism. Foundations of Physics 50, 1-26.

### v3

@@ 2.5 A survey of quantum probabilities
These results give an extended interpretation of the operator $A=A^{\theta}$: There is a simple formula for all expectations in terms of the operator. On the other hand, the set of such expectations determines the state of the system. Also on the other hand: If $A$ is specialized to the operator of an indicator function, we get back Born’s formula, so the consequences are equivalent to this formula.
A consequence of 3) above is that $\theta=\theta^{b}$ does not need to be maximal in order that a Born formula should be valid. Version 2) of Born’s formula is also valid when the basic variable $\theta^{a}$ behind the density operator $\rho$ is not maximal, under a certain assumption: Assume that $\theta^{a}=f(\lambda^{a})$, where $\lambda^{a}$ is maximal and the conditional probability distribution of $\lambda^{a}$, given $\theta^{a}$ is uniform.
Measurements of empirical variables are discussed in Helland (2021), but here we will look at the simpler case of a perfect measurement. Assume that we know the state $|\psi\rangle$ of a system and that we want to measure a new e-variable $\theta^{b}$. This can be discussed by means of the projection operators $\Pi_{j}^{b}=|b;j\rangle\langle b;j|$. First observe that by a simple calculation from Born’s formula
@@
It is interesting that Shrapnel et al. (2017), recently simultaneously derived both the Born rule and the well-known collapse rule from a knowledge-based perspective. The collapse rule is further discussed in Helland (2021), but in this article we will just assume this derivation as given. Then, after a perfect measurement $\theta^{b}=u_{j}^{b}$ has been obtained, the state changes to
@@ 2.6 Born’s rule in a decision setting
Go back to the person $C$ who is facing at least two related maximal decisions, with corresponding decision variables $\theta$ and $\eta$. Each of these decisions may be composed of several partial decisions; then the decision variables may be thought of as vectors. However, we will make the simplifying assumption that each partial decision has a finite number of outcomes; then the total decision also has a finite number of outcomes, and without loss of generality, we can order these outcomes, say $\eta$, as $1,2,...,r$.
According to Kahnemam (2011), some of our decisions are slow, and sometimes, in fact in very many cases, we have to make fast decisions, relying on earlier experience, perhaps in a subconscious way, relying on some abstract ideals that we have in our minds. This is consistent with the model assumptions that we have made above in connection to Born’s rule. And the fact that such fast decisions may be modelled by quantum probabilities, is consistent with a lot of empirical findings, see Pothos and Busemeyer (2022).
Focus again on the person $C$ and his fast decisions. Depending upon the situation that he is in, he may just have made several irrelevant decisions, collected in the variable $\eta$, but at the same moment, he is trying to focus on the problem that he really is interested in, and the corresponding (maximal) variable is called $\theta$. There is not much lack of generality in assuming that these two decision processes have the same number of outcomes $r$; we can only add some abstract irrelevant outcomes to one of the decision variables.
According to my interpretation of quantum theory and of decision making, each of the two decision variables has an operator connected to it, and the eigenspaces of such an operator correspond to a concrete outcome of the decision process. The probabilities are calculated as in (3). The application of these theoretical considerations is best studied by an example.
@@ 3 Quantum decision theory in practice; an example
Consider a situation with one doctor and one patient. The doctor has the choice between two mutually exclusive medicines $A$ and $B$.
The doctor starts by asking a number of questions, and he obtains answers. Result: A state $|\psi\rangle$ for the patient. This is the starting point of our model of what is going on in the mind of the doctor during the decision process. The state is modelled by some vector in an underlying Hilbert space, connected to the mind of the doctor.
@@
However, from our point of view, the most interesting case is where there is some coupling between the two medicines. The doctor must make fast decisions. In his mind he may have a lot of other decisions that he has made or is going to make. As a model, these may be collected in a decision variable $\eta$. But now he focuses on the actual patient. The decisions that he is going to make on this patient, are modelled, and collected in the decision variable $\theta$. In the doctor’s mind at the moment in question, is also all the information that he possesses about the medicines $A$ and $B$. As a doctor, he has ideals learned through many years of education and experience; these ideals can be modelled in terms of some abstract perfectly rational being $D$.
We will propose a quantum model with Hilbert space $\mathcal{H}$. The projector upon the subspace indicating that $A$ helps is denoted by $\Pi_{A}$.
@@
To continue the model, consider two orthogonal spaces in $\mathcal{H}$: $V_{B}$: Indicating that $B$ helps. The projector upon this space is $\Pi_{B}$. And $V_{B}^{\mathrm{perp}}$: Indicating that $B$ does not help. The projector in this space is $\Gamma_{B}=I-\Pi_{B}$.
@@ 3.1 Some consequences
For a tentative illustration, see Figure 1.
@@
Finally, note that one may have
@@
3.2 Interpretations
All these probabilities may be thought of as describing what is going on in the mind of the doctor before he makes a decision. Similar qualitative conducts of probabilities in the minds of people have been observed empirically in many investigations; see references in Pothos and Busemeyer (2013). It is interesting in itself that quantum probabilities may be connected to our minds.
In epistemic processes (processes to try to achieve new knowledge) quantum theory is also connected to the variables in the mind of an observer, or variables shared by a group of observers. This is related to our interpretation of the state vectors. But of course this interpretation - as any interpretation - may be controversial.
@@ 4 Concluding remarks
Quantum decision theory seems to be applicable in many cases where decisions are to be made, primarily but not only, in the cases that Kahneman (2011) calls decisions of type 1, fast decisions. When a prior is available, an alternative is a Bayesian decision theory. This theory will follow the Kolmogorov rules, and apart from the existence of a prior, the theory does not require any special assumptions.
In this article I have concentrated on decisions made by a single person. however, the theory can be extended to joint decisions made by a group of communicating persons. Then these persons must have common ideals, as modelled by a perfectly rational higher being $D$. The ideals must be common in the setting determined by the relevant decision variables.
@@
As a basis for quantum decision theory as discussed here, lies the idea of two complementary decision variables at the same time in the mind of some person or in the joint minds of a group of communicating persons. I repeat that two different variables are complementary if they are maximal as accessible variables. Another essential assumption is the existence of ideals that are perceived as perfectly rational, modelled by a hypothetical ‘ideal higher being’.
The quantum probabilities may be thought of as relative to this ‘ideal higher being’, hence they must be looked upon as absolute probabilities, not conditioned in any way. This is confirmed by empirical findings in the quantum decision literature.
There are cases where powerful political leaders may be roughly modelled as such ‘ideal higher beings’, relative to the influence they have on other people and their decisions. A fundamental requirement should then be that the leaders obey certain basic humanitarian conditions, including respect for other people’s lives. Unfortunately, this goal is not always satisfied in practice.
@@
The assumptions may in certain situations be satisfied by many different pairs of decision variables, leading to mental difficulties in the first case and to partly complementary world views in the last two cases. Then in these cases it might imply deep conflicts between those persons or groups if some necessary conditions are there. We can see many examples of this phenomenon in the world today.
5 Aknowledgments